OpenStudy (anonymous):

If cosθ=−3/4 with π<θ<3π/2 , find the following: (a) cos (−θ) ,(b) sec(−θ) ,(c) sin(−θ) If cosθ=−3/4 with π<θ<3π/2 , find the following: (a) cos (−θ) ,(b) sec(−θ) ,(c) sin(−θ) @Mathematics

OpenStudy (anonymous):



OpenStudy (anonymous):

use the appropriate ratios, making sure to be careful with the signs. \[\cos(-\theta)=\cos(\theta)=-\frac{3}{4}\] because cosine is even

OpenStudy (anonymous):

help me do the sec and sine

OpenStudy (anonymous):

the sine becomes what?